OpenStudy (anonymous):

Which ordered triple is a solution of the following linear system?: 2x + 3y + 2z = 26 -x + 7y - z = 21 3x + 2y - 5z = 13

OpenStudy (freckles):

elimination is needed

OpenStudy (freckles):

can we do matrices?

OpenStudy (anonymous):

idk what that is

OpenStudy (freckles):

ok well again try multiply the second equation by 2 and adding the new equation to the first equation

OpenStudy (freckles):

you have first equation is 2x+3y+2z=26 second equation is -2x+14y-2z=42 ----------------- now add these equations together

OpenStudy (anonymous):

i got .. -2+14y-2z=42

OpenStudy (freckles):

2x+(-2x) is not -2 3y+14y is not 14y 2x+(-2x) is not -2z 26+42 is not 42

OpenStudy (anonymous):

im confused u lost me now .. u said to multiply the equation by 2 .

OpenStudy (freckles):

the second by positive 2 yes

OpenStudy (freckles):

you have first equation is 2x+3y+2z=26 second equation is -2x+14y-2z=42 ----------------- now add these equations together

OpenStudy (anonymous):

17y=68 y= 4

OpenStudy (freckles):

yes! i have to go sorry

OpenStudy (anonymous):

noooo . please juss 5 more mins.

OpenStudy (freckles):

i'm back if you are still here